# LangTime: A Language-Guided Unified Model for Time Series Forecasting with Proximal Policy Optimization

### v1

@@ Abstract
Large language models, Time series, Reinforcement learning
@@ 1 Introduction
Time series refers to sequences of data points indexed in discrete-time order (Box et al., 2015), and they are common in real-world applications, such as financial risk assessment (Baffour et al., 2019), energy sustainability (Yu et al., 2023), and weather forecasting (Yu et al., 2025; Sun et al., 2021). Recently, large language models (LLMs) (Radford, 2018) have demonstrated remarkable capabilities in capturing sequential structures and patterns, which is crucial for modeling time-dependent data in time series forecasting. Given the similarities between time series and natural language in sequence modeling, several approaches have successfully leveraged pre-trained language models for time series forecasting, yielding promising outcomes (Xue & Salim, 2023; Gruver et al., 2024; Pan et al., 2024; Jia et al., 2024).
@@
Figure 1: Challenges of Applying LLMs to Time Series Data. (a) Cross-domain generalization: Different domains possess their unique characteristics and numerical implications. (b) Cross-modality alignment: LLMs struggle to directly comprehend unseen time series as they are trained on language data, whereas language guidance can assist in enhancing their understanding. (c) Error accumulation: In autoregressive prediction, using the output from the previous step as input is considered unreliable.
However, as shown in Figure 1, arising from the inherent differences between time series and natural language, applying LLMs to time series forecasting presents three primary challenges. Firstly, cross-domain generalization. In contrast to natural language, which is governed by structural and semantic rules that ensure consistency across various domains, time series display a wide array of statistical characteristics across different fields. Furthermore, identical values may possess disparate meanings across distinct domains, thereby complicating the task of integrating multi-domain information into LLMs effectively. To address this challenge, some methods leverage semantic information embedded in datasets as prompts, enabling LLMs to effectively differentiate and adapt to various domains (Liu et al., 2024d). Secondly, cross-modality alignment. Pre-trained on an extensive text corpus, LLMs exhibit limited capacity for directly understanding time series, thereby rendering cross-modal alignment a formidable challenge (Zhou et al., 2023). While language as prefixes partially integrates temporal information, they fall short of achieving seamless alignment between the two modalities, limiting meaningful interaction and effective integration (Jin et al., 2023).
In addition, error accumulation in autoregressive frameworks. For tasks with varying prediction horizons, autoregressive models allow a single framework to handle multiple horizons, avoiding the need for separate training protocols (Liu et al., 2024f, e). However, supervised training is limited to optimizing a model’s ability to predict the next step based on actual historical observations. As a result, it fails to alleviate the adverse effects of error accumulation in autoregressive prediction.
@@ 2.1 Large Language Models for Time Series
Figure 2: Overview of LangTime: (1) LangTime constructs Temporal Comprehension Prompts to help LLMs understand <Time Series Representation>, which is generated by the Temporal Encoder. (2) Alignment of LLMs with Time series during the pre-training via compressed representation learning and reconstruction-prediction tasks. (3) TimePPO fine-tune algorithm to mitigate error accumulation and enhance long-term forecasting. The content of the prompt has been simplified in the figure, with the detailed information provided in Section 3.3 and Section A.2.
@@ 3.1 Architecture
Pre-training. To align time series representation with the word embedding space, we utilize two complementary pre-training tasks: reconstruction and prediction. The reconstruction task enhances the model’s understanding of time series, while the prediction task leverages the generative capabilities of LLMs to identify anticipatory dependencies for future forecasting (Cao et al., 2020). We adopt the Huber loss (Huber, 1992) to balance robustness and sensitivity. The overall pre-training loss is formulated as:
@@ 3.3 Temporal Comprehension Prompts
Time: <Time Series Representation>,
@@
All TCP components are fed into the pre-trained LLM, where the causal attention mechanism enables each token to incorporate information from all preceding tokens. Following (Zhang et al., 2024), we extract the previous token from <|EMB|>  and <|OUT|>  respectively, serving as the compressed token and prediction token. Ultimately, we employed a projection operation to obtain the reconstruction and prediction:
@@ 3.4 Time Series Proximal Policy Optimization
Figure 3: Overview of TimePPO. Both the Reference Model and the Fine-Tuned Model are initialized with the same parameters.
@@
$R(\hat{y}_{t})=$ $\text{tanh}\left(\tau\sum_{i}\mathcal{R}_{i}(\hat{y}_{t},y_{t})w_{i}\right)$ (5)
$-\beta MSE(\hat{y}_{t},\hat{y}_{t}^{\text{ref}}),$
@@
$+\eta\mathbb{E}||y-\hat{y}||_{2}^{2},$ (8)
where $r(\theta)$ represents the policy ratio, and $\epsilon$ regulates the extent of policy updates to ensure stability. The alignment tax term $\eta\mathbb{E}||y-\hat{y}||_{2}^{2}$ penalizes large deviations from the ground truth, ensuring that the model remains aligned with accurate predictions. Further details are presented in Section A.4.
@@
* indicates the adoption of the official baseline code, with adjustments to input sequence length and maximum training epochs for fair comparison with other methods, and other results are sourced from TimesNet (Wu et al., 2022).
@@ 4.1 Training Details
Datasets. We extensively evaluate LangTime on seven real-world benchmark datasets spanning various time series application domains, i.e., ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange, and Weather. The details of these datasets are shown in the Section B.1
Baselines. We evaluate LangTime against state-of-the-art models. (1) LLM-based methods: UniTime (Liu et al., 2024d), AutoTimes (Liu et al., 2024e), S2IP-LLM (Pan et al., 2024), Time-LLM (Jin et al., 2023), and GPT4TS (Zhou et al., 2023); (2) Specific models: PatchTST (Nie et al., 2022), and TimesNet (Wu et al., 2022). We adopt Qwen2-0.5B-Instruction (Yang et al., 2024) as our backbone. Detailed implementations are provided in Appendix B.
Setup. We adopt pre-training to fine-tuning dataset ratio of $\mathcal{D}_{\text{pre-train}}:\mathcal{D}_{\text{fine-tune}}=8:2$. To ensure a fair comparison, all methods maintained a fixed look-back window of $L=96$ and predicted future values with lengths of $F=\{96,192,336,720\}$.
@@ 4.2 Main Results
Pre-trained LangTime achieves remarkable improvements over baseline models that are also trained across datasets, achieving the best performance in 58 out of 70 entries. This highlights LangTime’s ability to generalize effectively across diverse time series distributions. On the right side of the table, results indicate that TimePPO fine-tuned LangTime achieves competitive performance, surpassing models trained individually on each dataset in 50 out of 70 entries, establishing new state-of-the-art results. These findings validate LangTime’s effectiveness in handling time series data with diverse characteristics. Furthermore, compared to other LLM-based forecasting methods, LangTime demonstrates significant advantages, reinforcing its capability to enhance the alignment between LLMs and time series data.
Comparison with Other Fine-tuning Methods. Table 2 presents a comparison between our proposed TimePPO algorithm and Supervised Fine-tuning (SFT). Across most datasets, models fine-tuned with TimePPO achieve superior predictive performance compared to those fine-tuned with SFT. This demonstrates the effectiveness of TimePPO in mitigating cumulative errors and enhancing the stability of autoregressive forecasting models.
@@
Model ETTh1 Weather
@@ 4.3 Ablation Studies
Rewards Function. We evaluate LangTime’s predictions across three different dimensions and compute the corresponding reward scores, with the results presented in Table 4. The findings indicate that incorporating all dimensions significantly enhances the performance of pre-trained LangTime. Conversely, removing any single dimension leads to a decline in TimePPO’s effectiveness, with Mean Squared Error (MSE) exerting the most significant impact compared to the other two dimensions.
@@
4.4 Zero-Shot Transferability Analysis
In this section, we conduct an in-depth analysis of the zero-shot performance of our pre-trained model compared to other cross-domain training models in unseen domains. Table 5 demonstrates that LangTime consistently outperforms baseline models in most cases. This highlights the effectiveness of TCPs in enhancing the model’s adaptability to diverse, unseen time series distributions.
@@ 4.5 Parameter Analysis
The Sensitivity of $\tau$. Figure 5 illustrates the impact of $\tau$ on TimePPO’s rewards function. Since $\tau$ influences the distribution of reward scores, its effect varies based on the dataset’s MSE and MAE convergence values. For ETTh1, performance remains relatively stable across different values of $\tau$, indicating low sensitivity. In contrast, for Weather, an improper $\tau$ selection leads to noticeable fluctuations, occasionally causing performance degradation after fine-tuning. Nevertheless, in most cases, $\tau$ enhances performance, demonstrating the robustness of our proposed rewards function.
@@ 4.6 Model Analysis
T-SNE Visualization. To mitigate channel confusion in training, we incorporate channel-wise instructions in TCPs. To further analyze its impact, we compare the mean pooling of temporal representations and compressed tokens generated by LLMs for the first three channels of the ETTh1 and Weather datasets using T-SNE visualization.
As shown in Figure 6, the TE processes all channels uniformly, resulting in overlapping representations across domains, particularly between High Useful Load and High Useless Load in ETTh1. However, with the guidance of TCPs, LLMs successfully differentiate these channels and capture their semantic relationships. For instance, Air Temperature and Potential Temperature exhibit stronger correlations compared to Air Pressure, reflecting intrinsic dependencies within the dataset.
@@
Figure 6: T-SNE visualization of the temporal representations and compressed token.
Figure 7: Attention map visualization for the last token. Special tokens and system prompts are hidden for a more intuitive display.
Attention Map Analysis. To further examine how LLMs interpret and integrate time series representations, we visualize the last token attention weights from the final layer of the LLM in LangTime, as shown in Figure 7. The results indicate that: ① The last token focuses primarily on the time series representations and the token preceding the compressed token placeholder, which plays a key role in reconstructing the input sequence. ② LLMs also attend to contextual and prediction-relevant information, reinforcing the effectiveness of TCPs in aligning time series with LLMs. These findings confirm that TCPs enhance LLMs’ ability to differentiate time series patterns, preserve domain-specific knowledge, and capture meaningful temporal dependencies.

### v2

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Challenges of Applying LLMs to Time Series. (a) Cross-domain generalization: Different domains possess their unique characteristics and numerical implications. (b) Cross-modality alignment: LLMs struggle to directly comprehend unseen time series as they are trained on language data, whereas language guidance can assist in enhancing their understanding. (c) Error accumulation: In autoregressive prediction, using the output from the previous step as input is considered unreliable.
Time series refers to sequences of data points indexed in discrete-time order (Box et al., 2015), and they are common in real-world applications, such as financial risk assessment (Baffour et al., 2019), energy sustainability (Yu et al., 2023b), and weather forecasting (Yu et al., 2025; Sun et al., 2021). The rapid development of machine learning has driven significant advances in time series forecasting(Yu et al., 2023a; Shao et al., 2024; Sun et al., 2025). Recently, large language models (LLMs) (Radford, 2018) have demonstrated remarkable capabilities in capturing sequential structures and patterns, which is crucial for modeling time-dependent data in time series forecasting. Given the similarities between time series and natural language in sequence modeling, several approaches have successfully leveraged pre-trained language models for time series forecasting, yielding promising outcomes (Xue & Salim, 2023; Gruver et al., 2024; Pan et al., 2024; Jia et al., 2024).
However, as shown in Figure 1, arising from the inherent differences between time series and natural language, applying LLMs to time series forecasting presents three primary challenges. Firstly, cross-domain generalization. Unlike natural language with domain-consistent structural/semantic rules, time series exhibit diverse statistical patterns across domains. Moreover, identical values may carry domain-specific meanings, hindering effective multi-domain integration into LLMs. To address this challenge, some methods leverage semantic information embedded in datasets as prompts, enabling LLMs to effectively differentiate and adapt to various domains (Liu et al., 2024d). Secondly, cross-modality alignment. Pre-trained on an extensive text corpus, LLMs exhibit limited capacity for directly understanding time series, thereby rendering cross-modal alignment a formidable challenge (Zhou et al., 2023). While language-as-prefixes models concatenate the two modalities, the lack of meaningful interaction hinders seamless alignment (Jin et al., 2023).
In addition, error accumulation in autoregressive frameworks. For tasks with varying prediction horizons, autoregressive models allow a single framework to handle multiple horizons, avoiding the need for separate training protocols (Liu et al., 2024f, e; Yu et al., 2024). However, supervised training is limited to optimizing a model’s ability to predict the next step based on actual historical observations. As a result, it fails to alleviate the adverse effects of error accumulation in autoregressive prediction.
@@ 2.1 Large Language Models for Time Series
Figure 2: Overview of LangTime: (1) The Temporal Encoder extracts temporal dependencies from time series data of varying lengths and domains, generating unified time series representations. (2) Temporal Comprehension Prompts are constructed by integrating time series representations with domain-specific context, enabling LLMs to perform temporal pattern compression. This pre-training phase establishes modality alignment by jointly optimizing reconstruction and prediction tasks. (3) During fine-tuning, the TimePPO algorithm is implemented to mitigate error accumulation in autoregressive forecasting, thereby enhancing long-term prediction. The content of the prompt has been simplified in the figure, with the detailed information provided in Section 3.3 and Section A.2.
@@ 3.1 Architecture
Multi-task Pre-training. To align time series representation with the word embedding space, we utilize two pre-training tasks: reconstruction and prediction. The reconstruction task enhances the model’s understanding of time series, while the prediction task leverages the generative capabilities of LLMs to identify anticipatory dependencies for future forecasting (Cao et al., 2020). We adopt the Huber loss (Huber, 1992) to balance robustness and sensitivity. The overall pre-training loss is formulated as:
@@ 3.3 Temporal Comprehension Prompts
Value: <Time Series Representation>,
@@
All components of the TCPs are fed into the pre-trained LLM, where the causal attention mechanism enables each token to incorporate information from all preceding tokens. Following (Zhang et al., 2024), we extract the previous token from <|EMB|>  and <|OUT|>  respectively, serving as the compressed token and prediction token. Ultimately, we employed a projection operation to obtain the reconstruction and prediction:
@@ 3.4 Time Series Proximal Policy Optimization
Figure 3: Overview of TimePPO. Both the Reference Model and the Fine-Tuned Model are initialized with the same parameters. The Value Function uses a repeat strategy to estimate expected returns, while the Rewards Function evaluates the accuracy of prediction results through multiple dimensions.
@@
$R(\hat{y}_{t})=\text{tanh}\left(\tau\sum_{i}\mathcal{R}_{i}(\hat{y}_{t},y_{t})w_{i}\right)-\beta MSE(\hat{y}_{t},\hat{y}_{t}^{\text{ref}}),$ (5)
@@
$-\eta\mathcal{L}_{\text{prediction}},$ (8)
where $r(\theta)$ represents the policy ratio, and $\epsilon$ regulates the extent of policy updates to ensure stability. The alignment tax term $\eta\mathcal{L}_{\text{prediction}}$ penalizes large deviations from the ground truth, ensuring that the model remains aligned with accurate predictions. Further details are presented in Section A.4.
@@
* indicates the adoption of the official baseline code, with adjustments to input sequence length and maximum training epochs for fair comparison with other methods, and other results are sourced from iTransformer(Liu et al., 2023).
@@ 4.1 Training Details
Datasets. We conduct both pre-training and fine-tuning of LangTime on seven real-world datasets that cover various time series application domains, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange, and Weather. Specifically, we pre-train LangTime across multiple domains and then perform fine-tuning on individual domain using the TimePPO algorithm. We evaluate the performance of LangTime on the test splits of the same seven benchmark datasets. The details of these datasets are provided in Section B.1.
Baselines. We evaluate LangTime against state-of-the-art models. (1) LLM-based methods: UniTime (Liu et al., 2024d), AutoTimes (Liu et al., 2024e), S2IP-LLM (Pan et al., 2024), Time-LLM (Jin et al., 2023), and GPT4TS (Zhou et al., 2023); (2) Specific methods: PatchTST (Nie et al., 2022), and TimesNet (Wu et al., 2022). We adopt Qwen2-0.5B-Instruction (Yang et al., 2024) as backbone.
Setup. We adopt pre-training to fine-tuning dataset ratio of $\mathcal{D}_{\text{pre-train}}:\mathcal{D}_{\text{fine-tune}}=8:2$. To ensure a fair comparison, all methods maintained a fixed look-back window of $L=96$ and predicted future values with lengths of $F=\{96,192,336,720\}$. More implementation details can be found in Section B.2.
@@ 4.2 Main Results
Pre-trained LangTime achieves remarkable improvements over baseline models that are also trained across datasets, achieving the best performance in 58 out of 70 entries. This highlights LangTime’s ability to generalize effectively across diverse time series distributions. On the right side of the table, results indicate that TimePPO fine-tuned LangTime achieves competitive performance, surpassing models trained individually on each dataset in 45 out of 70 entries, establishing new state-of-the-art results. These findings validate LangTime’s effectiveness in handling time series data with diverse characteristics. Furthermore, compared to other LLM-based forecasting methods, LangTime demonstrates significant advantages, reinforcing its capability to enhance the alignment between LLMs and time series data.
Comparison with Other Fine-tuning Algorithms. Table 2 presents a comparison between our proposed TimePPO algorithm and Supervised Fine-tuning algorithm (SFT). Across most datasets, models fine-tuned with TimePPO achieve superior predictive performance compared to those fine-tuned with SFT. This demonstrates the effectiveness of TimePPO in mitigating cumulative errors and enhancing the stability of autoregressive forecasting models.
@@
Method ETTh1 Weather
@@ 4.3 Ablation Studies
Rewards Function. We comprehensively evaluate the accuracy of prediction results from three dimensions in the TimePPO, including Mean Squared Error (MSE), Mean Absolute Error (MAE), and KL divergence, as detailed in Table 10. We evaluate LangTime’s predictions across the three different dimensions and compute the corresponding reward scores, with the results presented in Table 4. The findings indicate that incorporating all dimensions significantly enhances the performance of pre-trained LangTime. Conversely, removing any single dimension leads to a decline in TimePPO’s effectiveness, with MSE exerting the most significant impact compared to others.
Backbones. We selected Qwen2-0.5B-Instruction (Yang et al., 2024) as the backbone due to its superior instruction-following capability. To validate the framework’s scalability, we substituted the backbone with GPT-2 (Radford et al., 2019). Table 5 demonstrates that while GPT-2 achieves comparable performance on the Weather dataset, it underperforms on ETTh1 due to weaker instruction compliance. To evaluate LLMs’ role in transferring sequential modeling from text to time series, we replaced the backbone with a linear layer (Zeng et al., 2023). As presented in Table 5, this modification resulted in significant performance degradation, conclusively demonstrating that our method effectively enhances LLMs’ comprehension of time series patterns.
Table 5: Ablation study of different backbones on ETTh1 and Weather datasets. Bold: best results.
Dataset GPT2 Qwen Linear
MSE MAE MSE MAE MSE MAE
ETTh1 96 0.409 0.406 0.388 0.391 0.702 0.567
192 0.472 0.438 0.442 0.423 0.721 0.580
336 0.530 0.461 0.479 0.445 0.733 0.592
720 0.556 0.483 0.482 0.465 0.735 0.611
Avg 0.492 0.447 0.448 0.431 0.723 0.588
Weather 96 0.194 0.233 0.181 0.217 0.223 0.273
192 0.244 0.276 0.236 0.254 0.263 0.310
336 0.294 0.311 0.294 0.295 0.318 0.343
720 0.366 0.358 0.368 0.341 0.398 0.405
Avg 0.275 0.294 0.270 0.277 0.301 0.333
Adaptability to Different Prompts. The TCPs in our approach comprise two components. The background part incorporates domain and channel descriptions to provide richer linguistic information, enabling LLMs to develop a more profound understanding of time series background knowledge. The instruction part has two task instructions to guide the model in understanding the time series data and generating predictions. We compared the impact on performance when these two parts varied separately, as shown in Table 6. However, when their expression changes but the basic meaning remains consistent, modifying the domain description and instructions does not significantly affect the model’s performance. The detailed modification of the prompt can be found in Section B.3.
Table 6: Impact of language prompt modification on model capability on ETTh1 and Weather datasets. Bold: best performance.
Dataset Original Instruction Background
MSE MAE MSE MAE MSE MAE
ETTh1 96 0.388 0.391 0.386 0.397 0.387 0.396
192 0.442 0.423 0.442 0.425 0.441 0.426
336 0.479 0.445 0.479 0.445 0.478 0.445
720 0.482 0.465 0.488 0.465 0.493 0.468
Avg 0.448 0.431 0.449 0.433 0.450 0.434
Weather 96 0.181 0.217 0.181 0.219 0.184 0.214
192 0.236 0.254 0.241 0.261 0.237 0.256
336 0.294 0.295 0.297 0.300 0.296 0.298
720 0.368 0.341 0.371 0.347 0.372 0.348
Avg 0.270 0.277 0.273 0.282 0.272 0.279
4.4 Zero-Shot Transferability Analysis
@@
In this section, we conduct an in-depth analysis of the zero-shot performance of our pre-trained model compared to other cross-domain training models in unseen domains. Table 7 demonstrates that LangTime consistently outperforms baseline models in most cases. This highlights the effectiveness of TCPs in enhancing the model’s adaptability to diverse, unseen time series distributions. Additionally, we further analyze LangTime’s cross-domain transferability in Section C.1.
@@ 4.5 Parameter Analysis
The Sensitivity of $\tau$. Figure 5 illustrates the impact of $\tau$ on TimePPO’s rewards function. Since $\tau$ influences the distribution of reward scores, its effect varies based on the dataset’s MSE and MAE convergence values. For ETTh1, performance remains relatively stable across different values of $\tau$, indicating low sensitivity. In contrast, for Weather, an improper $\tau$ selection leads to noticeable fluctuations, occasionally causing performance degradation after fine-tuning. Nevertheless, in most cases, TimePPO exhibits robustness with respect to $\tau$ within a certain range.
The Sensitivity of $\beta$ and $\eta$. Figure 6 illustrates the model’s sensitivity to the parameters $\beta$ and $\eta$. Here, $\beta$ controls the MSE penalty term, aiming to regulate the extent of policy updates from the reward score perspective. For $\eta$, we referenced the alignment tax design in InstructGPT (Ouyang et al., 2022), enhancing the model’s prediction capability and training stability. The sensitivity analysis experiments revealed no significant impact of these parameters, demonstrating the robustness of our proposed method.
(a) $\beta$
(b) $\eta$
Figure 6: Parameter sensitivity analysis on $\beta$ and $\eta$.
@@ 4.6 Model Analysis
Error Accumulation Analysis. To evaluate how TimePPO improves model resilience against cumulative errors, we analyzed prediction metrics at the final step of long-term forecasting (336 and 720). As shown in Table 8, TimePPO demonstrates superior performance in these critical final-step predictions where accumulated errors have greatest impact. This improvement stems from TimePPO’s unique approach: its value function estimates sequence-wide returns while advantage calculations assess each step’s long-term value relative to these estimates. Unlike methods that focus solely on immediate next-step prediction, these mechanisms optimize the entire sequence’s prediction quality. Through this holistic optimization, TimePPO effectively mitigates the cumulative error problem inherent in autoregressive models.
Table 8: Comparison of algorithms in mitigating accumulated error on ETTh1 and Weather datasets. Reporting last-step metrics for long-term autoregressive forecasting (336, 720), with last 96 points for 336 and 48 points for 720.
Dataset Pre-Training SFT TimePPO
MSE MAE MSE MAE MSE MAE
ETTh1 336 0.527 0.469 0.526 0.468 0.519 0.464
720 0.541 0.513 0.539 0.512 0.537 0.513
Avg 0.534 0.491 0.533 0.490 0.528 0.489
Weather 336 0.364 0.353 0.364 0.353 0.358 0.350
720 0.467 0.418 0.475 0.418 0.462 0.412
Avg 0.415 0.385 0.419 0.386 0.410 0.381
T-SNE Visualization. In this part, we focus on analyzing the three channels of the ETTh1 and Weather datasets. The ETTh1 dataset emphasizes energy-related features, including High Useful Load, High Useless Load, and Middle Useful Load, while the Weather dataset comprises features related to the natural environment, such as Air Temperature, Air Pressure, and Potential Temperature. We employ T-SNE visualization techniques to analyze LangTime’s performance in processing these features, aiming to better understand the model’s adaptability across different domains.
As shown in Figure 7, initially, TE processed all channels uniformly, leading to overlaps in representations between different domains such as High Useful Load and High Useless Load within the ETTh1 dataset. However, TCPs provided channel guidance to LLMs, enabling them to effectively distinguish these features and uncover their semantic relationships. Specifically, in the Weather dataset, Air Temperature and Potential Temperature exhibit stronger correlation compared to Air Pressure, demonstrating the role of language guidance in uncovering intrinsic dependencies between different channels.
@@
Figure 7: T-SNE visualization of the mean pooling of temporal representations generated by Temporal Encoder and compressed token generated by LLM.
Figure 8: Attention map visualization for the last token, which is the actual prediction token preceding <|OUT|>. Special tokens and system prompts are hidden, and softmax values are recalculated over the remaining tokens to provide a clearer representation.
Attention Map Analysis. To further examine how LLMs interpret and integrate time series representations, we visualize the last token attention weights from the final layer of the LLM in LangTime that is ultimately processed through the MLP to generate predictions, as shown in Figure 8. The results indicate that: ① The last token focuses primarily on the time series representations and the token preceding the compressed token placeholder, which plays a key role in reconstructing the input sequence. ② LLMs also attend to contextual and prediction-relevant information, reinforcing the effectiveness of TCPs in aligning time series with LLMs. These findings confirm that TCPs enhance LLMs’ ability to differentiate time series patterns, preserve domain-specific knowledge, and capture meaningful temporal dependencies.